OpenStudy (anonymous):

Find the 10th term of the sequence whose nth term is given by the formula tn= (-1)^n+1/n

OpenStudy (anonymous):

where you see an \(n\) replace it by a \(10\)

OpenStudy (anonymous):

so is it -1/10?

OpenStudy (anonymous):

not the way i read it no

OpenStudy (anonymous):

is it \[t_n=(-1)^n+\frac{1}{n}\] or \[t_n=(-1)^n\times \frac{1}{n}\]?

OpenStudy (anonymous):

no its tn (-1) ^ n+1/n

OpenStudy (anonymous):

its either -1/10, 1/10, -1/7 or 1/7

OpenStudy (anonymous):

\[(-1)^{10}+\frac{1}{10}=1+\frac{1}{10}=\frac{11}{10}\] so perhaps it is not the right \(t_n\)

OpenStudy (anonymous):

unless you mean \[\frac{(-1)^{n+1}}{n}\] could that be it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats what I meant to put

OpenStudy (anonymous):

then compute \[\frac{(-1)^{10+1}}{10}\]

OpenStudy (anonymous):

thats what I thought thanks :)

OpenStudy (anonymous):

yw